# PKU-DyMVHumans: A Multi-View Video Benchmark for High-Fidelity Dynamic Human Modeling

### v1

@@ 5 Conclusion
In this work, we introduced PKU-DyMVHumans, a dynamic human dataset designed for high-fidelity human reconstruction and rendering from dense multi-view videos. It features high resolution, cloth deformation, complex motion, intricate texture details, and human-object interactions, as well as challenging multi-person interactions and complex scene effects (e.g., lighting, shadows, and smoking). We further presented benchmark tasks, with detailed experiments on several advanced methods. PKU-DyMVHumans further fill in the gap between existing datasets and real-scene applications.
Challenges and future works. While we have validated the complexity and fidelity of our dataset on numerous human-centric reconstruction and rendering. It is significant to highlight the more challenging and realistic multiple person/subject modeling that could reflect the rendering differences with respect to multi-persons interactivity, complex scene effects, and multi-view consistent performance. Additionally, free-viewpoint rendering of a moving subject from a monocular self-rotating video is a complex yet desirable setup. Our supplementary material provides additional experiments for free-viewpoint rendering of moving subject, the results are affected by local occlusion and view absence, leading to artifacts in view rendering. With these opportunities and challenges, we believe PKU-DyMVHumans will benefit the development of new approaches in the community.
@@
  • More dataset information in Sec. A, including more visualizations of data samples, and per-category data distribution;
@@ B.1 Neural Scene Decomposition
Method overview. For the scene captured by multi-view images, we use COLMAP [34] and RVM [20] to get sparse 3D points and coarse object masks as co-inputs, and predict a dense, geometrical consistent object map, as well as a textural, completed background for each image. We present a self-supervised learning framework towards delicate segmentation by combining the neural surface representation power from multi-view image observations via volume rendering of SDF. Fig. 12 shows an overview of our method, in which multi-view geometric constraints are embedded in the form of dense one-to-one mapping in 3D surface representation. By connecting the SDF-based surface representation to geometric consistency, and applying volume rendering to train this representation with robustness, we can reconstruct the foreground object geometry and appearance over time.
Figure 12: We present a self-supervised learning framework towards delicate segmentation by combining 3D neural surface representation power from multi-view images of a scene.
Network Architecture. To tackle this challenging task by leveraging the existence of geometric consistency of the one-to-one dense mapping in 3D space, we decouple the scene into two complementary neural representation modules based on hash-encoded SDF: a Foreground Consistent Representation (FoCoR) module and a Background Completion (BaCo) module. To accelerate the training process, we incorporate SDF volumetric rendering with multi-resolution hash encodings [25]. As shown in Fig. 13, the network architecture consists of the following components: (a) two multi-resolution hash grids with 16 levels of different resolutions ranging from 16 to 2048; (b) two SDF networks modeled by a 1-layer MLP with 64 hidden units; (c) an RGB network modeled by a 2-layer MLP with 64 hidden units for consistent foreground object; (d) an RGB network modeled by a 2-layer MLP with 64 hidden units for static background.
Figure 13: A visualization of the architecture of FoCoR and BaCo module.
Implementation details. To train our system, we introduce photometric and geometric losses to supervise the training process, with the multi-view images serving as the primary supervision signal. Our objective is to achieve fine-grained object segmentation and analyze the correlation between the neural surface representation and object segmentation. In this study, we evaluate two approaches: NeRF-based segmentation, which does not introduce the normal to regularize the output SDF implicitly, and SDF-based segmentation, which provides the SDF-based surface representation for cross-view geometric constraints.
3D Segmentation of scenes with a single/multiple foreground object. As shown in Fig. 14, our method successfully refines the segmentation remarkably using two neural representation models. When the normal is not introduced to implicitly regularize the output SDF (i.e., NeRF-based segmentation), it often produces noisy segmentation. However, when providing the SDF-based surface representation, the network is able to learn 3D geometry implicitly and generate an accurate foreground decomposition. These examples demonstrate that accurate prediction of object geometry with SDF-based surface representation is beneficial for object segmentation. However, when dealing with scenes involving multiple subjects, such as human-human and human-object interactions, as well as occlusion, it faces even greater challenges. Applying more advanced research to our dataset in the future will undoubtedly enhance the effects.
@@ B.2 Novel View Synthesis
Dance
1080_Dance_Dunhuang_Pair_f14f15
21.38 0.959 0.042 26.37 0.974 0.035 25.34 0.967 0.044
1080_Dance_Dunhuang_Single_f12
25.19 0.978 0.024 31.46 0.983 0.019 30.61 0.985 0.020
1080_Dance_Dunhuang_Single_f13
25.37 0.979 0.021 33.91 0.989 0.013 31.31 0.984 0.016
1080_Dance_Dunhuang_Single_f14
25.03 0.977 0.025 32.02 0.988 0.016 30.34 0.985 0.016
1080_Dance_Dunhuang_Single_f15
25.61 0.978 0.029 34.91 0.992 0.014 32.27 0.990 0.016
1080_Dance_Jazz_Single_c22
26.59 0.984 0.018 31.41 0.987 0.011 35.38 0.991 0.010
1080_Dance_Tibetan_Single_c22
26.26 0.984 0.020 33.69 0.991 0.016 34.28 0.990 0.014
1080_Dance_Banquet_Single_c23
26.73 0.984 0.016 36.20 0.993 0.009 35.56 0.993 0.009
@@
Kungfu
1080_Kungfu_Weapon_Pair_m12m13
19.51 0.941 0.061 22.31 0.955 0.014 22.07 0.962 0.048
1080_Kungfu_Double_Pair_m12m13
18.48 0.939 0.062 20.42 0.948 0.144 22.45 0.965 0.047
1080_Kungfu_Basic_Pair_c24c25
25.48 0.979 0.024 31.96 0.989 0.017 30.99 0.986 0.019
1080_Kungfu_Fan_Single_m12
25.72 0.981 0.024 33.16 0.988 0.021 30.56 0.985 0.021
1080_Kungfu_Taichi_Single_m12
22.16 0.976 0.027 30.94 0.988 0.020 29.60 0.986 0.020
1080_Kungfu_Shaolin_Single_m12
23.01 0.978 0.029 31.85 0.989 0.017 32.08 0.989 0.016
1080_Kungfu_Sword_Single_m13
23.24 0.978 0.023 31.10 0.986 0.019 29.39 0.985 0.019
1080_Kungfu_Spear_Single_m13
22.53 0.973 0.033 31.85 0.989 0.018 28.16 0.985 0.021
1080_Kungfu_Kick_Single_m13
20.16 0.970 0.032 29.43 0.986 0.032 28.68 0.987 0.024
1080_Kungfu_Basic_Single_m13
23.17 0.978 0.021 28.18 0.982 0.024 28.53 0.985 0.019
1080_Kungfu_Tongbeiquan_Single_m13
23.59 0.980 0.024 32.19 0.991 0.016 31.61 0.988 0.017
1080_Kungfu_Nunchuck_Single_m14
21.21 0.976 0.024 29.62 0.985 0.046 28.32 0.984 0.022
1080_Kungfu_Nanquan_Single_c24
25.81 0.983 0.026 37.62 0.995 0.013 34.67 0.993 0.014
1080_Kungfu_Broadsword_Single_c24
25.65 0.985 0.018 35.28 0.993 0.014 37.24 0.994 0.009
1080_Kungfu_Boxing_Single_c25
24.31 0.981 0.024 38.37 0.996 0.009 36.91 0.995 0.009
@@
Sport
1080_Sport_Football_Single_m11
24.91 0.983 0.017 29.83 0.982 0.018 30.50 0.986 0.016
1080_Sport_Taekwondo1_Pair_m11c21
23.72 0.970 0.037 32.50 0.989 0.019 27.12 0.981 0.029
1080_Sport_Badminton_Single_f11
25.22 0.980 0.028 34.29 0.993 0.011 33.79 0.993 0.014
@@
Fashion Show
4K_Studios_Show_Pair_f16f17
23.26 0.977 0.036 35.02 0.991 0.020 32.12 0.987 0.025
4K_Studios_Show_Pair_f18f19
22.80 0.976 0.031 34.24 0.993 0.016 32.23 0.992 0.014
4K_Studios_Show_Single_f16
20.38 0.975 0.042 34.49 0.990 0.012 34.33 0.993 0.012
4K_Studios_Show_Single_f17
23.07 0.982 0.036 35.08 0.992 0.021 34.11 0.992 0.018
4K_Studios_Show_Single_f18
22.95 0.983 0.027 36.73 0.995 0.012 34.32 0.994 0.013
4K_Studios_Show_Single_f19
24.36 0.986 0.024 38.94 0.996 0.010 37.03 0.995 0.011
4K_Studios_Dance_Single_f20
22.67 0.981 0.028 30.50 0.986 0.026 31.67 0.989 0.018

### v3

@@ 5 Conclusion
In this work, we introduced PKU-DyMVHumans, a dynamic human dataset designed for high-fidelity human reconstruction and rendering from dense multi-view videos. It features high-fidelity human performance, including high-detailed appearance, complex human motion, as well as challenging human-object interactions, multi-person interactions and complex scene effects (e.g., lighting, shadows, and smoking). We further presented benchmark tasks, with detailed experiments on several advanced methods. PKU-DyMVHumans further fill in the gap between existing datasets and real-scene applications.
Challenges and future works. While we have validated the complexity and fidelity of our dataset on numerous human-centric reconstruction and rendering scenarios. It is significant to highlight the more challenging and realistic multiple-person/subject modeling that could reflect the rendering differences with respect to multi-person interactivity, complex scene effects, and multi-view consistent performance. Additionally, free-viewpoint rendering of a moving subject from a monocular self-rotating video is a complex yet desirable setup. Our supplementary material provides additional experiments for free-viewpoint rendering of moving subjects, the results are affected by local occlusion and view absence, leading to artifacts in view rendering. With these opportunities and challenges, we believe PKU-DyMVHumans will benefit the development of new approaches in the community.
@@
  • More dataset information in Sec. A, including visualizations of data samples, and per-category data distribution;
@@ B.1 Neural Scene Decomposition
Method overview. For the scene captured by multi-view images, we use COLMAP [33] and BGMv2 [19] to get sparse 3D points and coarse object masks as co-inputs, and predict a dense, geometrical consistent object map, as well as a textural, completed background for each image. Fig. 12 shows an overview of our contemporary work Surface-SOS [53], in which multi-view geometric constraints are embedded in the form of dense one-to-one mapping in 3D surface representation. By connecting SDF-based surface representation to geometric consistency, and applying volume rendering to train the network with robustness, it can reconstruct the foreground object geometry and appearance over time.
Figure 12: Surface-SOS [53], a self-supervised learning framework towards delicate segmentation by combining 3D neural surface representation power from multi-view images of a scene.
Implementation details. To train Surface-SOS [53], we introduce photometric and geometric losses to supervise the training process, with the multi-view images serving as the primary supervision signal. Our objective is to achieve fine-grained object segmentation and analyze the correlation between the neural surface representation and object segmentation. In this study, we evaluate two approaches: NeRF-based segmentation, which does not introduce the normal to regularize the output SDF implicitly, and SDF-based segmentation, which provides the SDF-based surface representation for cross-view geometric constraints.
3D Segmentation of scenes with a single/multiple foreground object. As shown in Fig. 13, Surface-SOS successfully refines the segmentation remarkably using two neural representation models. When the normal is not introduced to implicitly regularize the output SDF (i.e., NeRF-based segmentation), it often produces noisy segmentation. However, when providing the SDF-based surface representation, the network is able to learn 3D geometry implicitly and generate an accurate foreground decomposition. These examples demonstrate that accurate prediction of object geometry with SDF-based surface representation is beneficial for object segmentation.
@@ B.2 Novel View Synthesis
Dance 1080_Dance_Dunhuang_Pair_f14f15 21.38 0.959 0.042 26.37 0.974 0.035 25.34 0.967 0.044
1080_Dance_Dunhuang_Single_f12 25.19 0.978 0.024 31.46 0.983 0.019 30.61 0.985 0.020
1080_Dance_Dunhuang_Single_f13 25.37 0.979 0.021 33.91 0.989 0.013 31.31 0.984 0.016
1080_Dance_Dunhuang_Single_f14 25.03 0.977 0.025 32.02 0.988 0.016 30.34 0.985 0.016
1080_Dance_Dunhuang_Single_f15 25.61 0.978 0.029 34.91 0.992 0.014 32.27 0.990 0.016
1080_Dance_Jazz_Single_c22 26.59 0.984 0.018 31.41 0.987 0.011 35.38 0.991 0.010
1080_Dance_Tibetan_Single_c22 26.26 0.984 0.020 33.69 0.991 0.016 34.28 0.990 0.014
1080_Dance_Banquet_Single_c23 26.73 0.984 0.016 36.20 0.993 0.009 35.56 0.993 0.009
@@
Kungfu 1080_Kungfu_Weapon_Pair_m12m13 19.51 0.941 0.061 22.31 0.955 0.014 22.07 0.962 0.048
1080_Kungfu_Double_Pair_m12m13 18.48 0.939 0.062 20.42 0.948 0.144 22.45 0.965 0.047
1080_Kungfu_Basic_Pair_c24c25 25.48 0.979 0.024 31.96 0.989 0.017 30.99 0.986 0.019
1080_Kungfu_Fan_Single_m12 25.72 0.981 0.024 33.16 0.988 0.021 30.56 0.985 0.021
1080_Kungfu_Taichi_Single_m12 22.16 0.976 0.027 30.94 0.988 0.020 29.60 0.986 0.020
1080_Kungfu_Shaolin_Single_m12 23.01 0.978 0.029 31.85 0.989 0.017 32.08 0.989 0.016
1080_Kungfu_Sword_Single_m13 23.24 0.978 0.023 31.10 0.986 0.019 29.39 0.985 0.019
1080_Kungfu_Spear_Single_m13 22.53 0.973 0.033 31.85 0.989 0.018 28.16 0.985 0.021
1080_Kungfu_Kick_Single_m13 20.16 0.970 0.032 29.43 0.986 0.032 28.68 0.987 0.024
1080_Kungfu_Basic_Single_m13 23.17 0.978 0.021 28.18 0.982 0.024 28.53 0.985 0.019
1080_Kungfu_Tongbeiquan_Single_m13 23.59 0.980 0.024 32.19 0.991 0.016 31.61 0.988 0.017
1080_Kungfu_Nunchuck_Single_m14 21.21 0.976 0.024 29.62 0.985 0.046 28.32 0.984 0.022
1080_Kungfu_Nanquan_Single_c24 25.81 0.983 0.026 37.62 0.995 0.013 34.67 0.993 0.014
1080_Kungfu_Broadsword_Single_c24 25.65 0.985 0.018 35.28 0.993 0.014 37.24 0.994 0.009
1080_Kungfu_Boxing_Single_c25 24.31 0.981 0.024 38.37 0.996 0.009 36.91 0.995 0.009
@@
Sport 1080_Sport_Football_Single_m11 24.91 0.983 0.017 29.83 0.982 0.018 30.50 0.986 0.016
1080_Sport_Taekwondo1_Pair_m11c21 23.72 0.970 0.037 32.50 0.989 0.019 27.12 0.981 0.029
1080_Sport_Badminton_Single_f11 25.22 0.980 0.028 34.29 0.993 0.011 33.79 0.993 0.014
@@
Fashion Show 4K_Studios_Show_Pair_f16f17 23.26 0.977 0.036 35.02 0.991 0.020 32.12 0.987 0.025
4K_Studios_Show_Pair_f18f19 22.80 0.976 0.031 34.24 0.993 0.016 32.23 0.992 0.014
4K_Studios_Show_Single_f16 20.38 0.975 0.042 34.49 0.990 0.012 34.33 0.993 0.012
4K_Studios_Show_Single_f17 23.07 0.982 0.036 35.08 0.992 0.021 34.11 0.992 0.018
4K_Studios_Show_Single_f18 22.95 0.983 0.027 36.73 0.995 0.012 34.32 0.994 0.013
4K_Studios_Show_Single_f19 24.36 0.986 0.024 38.94 0.996 0.010 37.03 0.995 0.011
4K_Studios_Dance_Single_f20 22.67 0.981 0.028 30.50 0.986 0.026 31.67 0.989 0.018